OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x^2 - 7x - 9. Find f(g(-1)). Select one: a. -21 b. -3 c. 3 d. 21

OpenStudy (johnweldon1993):

So you deal with the inside function first...g(x) ..when you plug -1 in for the 'x' in g..what does the equation come out to?

OpenStudy (anonymous):

I literally have no idea how to do this, I literally don't understand at all so can you tell me the first step

OpenStudy (johnweldon1993):

Sure! So first step...is look at the inside function...in f(g(-1)) this is g(x)...so look at your 'g' equation g(x) = x^2 - 7x - 9 when it has g(-1) it's telling you to replace every 'x' you see in the equation with that number (-1) ...so g(-1) = (-1)^2 - 7(-1) - 9 g(-1) = 1 + 7 - 9 g(-1) = ?

OpenStudy (anonymous):

If I plug in -1 for g(x) it's -1^2-7(-1)-9 = 1 + 7 - 9 = -1

OpenStudy (anonymous):

So -1

OpenStudy (johnweldon1993):

Right...now what you do...is you take what you got (-1) and you plug that into your OUTSIDE function now...here...that would be f(x) f(x) = x - 2 Plug in the -1 you just got for every 'x' you see.... f(-1) = -1 - 2 f(-1) = ?

OpenStudy (anonymous):

-3

OpenStudy (johnweldon1993):

And that would be correct :)